OpenStudy (anonymous):

how does 4^log2(x) equal to x^2?

OpenStudy (anonymous):

2 is the base??

OpenStudy (anonymous):

yeah

OpenStudy (kinggeorge):

\[\large 4^{\log_2(x)}=2^{2^{\log_2(x)}}=2^x\]not \(x^2\)

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

but wolfram alpha says the alternative form is x^2 http://www.wolframalpha.com/input/?i=4%5Elog2x

OpenStudy (lgbasallote):

@KingGeorge made a mistake... \[\Large 2^{2\log_2 x} = 2^{\log_2 x^2} = x^2\]

OpenStudy (kinggeorge):

That would appear to be correct. However, you should be careful about switching forms like that because \(\log_2(x)\) is only defined for positive numbers.

OpenStudy (anonymous):

oh I see Thanks!